# DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning

### v1

@@ 3.1. Multi-Objective Partition Allocation
where $x$ denotes a solution in the decision space $\chi\subset\mathbb{R}^{n}$, $\mathbf{f}:\chi\to\mathbb{R}^{m}$ represents an $m$-dimensional vector-valued objective function, and $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k}\,\text{where}\,k=1,...,n$ stands for a series of soft constraints. Importantly, the evaluation process is computationally expensive for each individual objective $f_{i}\left(x\right),i=1,...,m$.
@@
Setup.  Suppose that we have a training corpus $\mathcal{D}$ and training sequences $\mathcal{W}$ pending sampling, the context window for each is limited in length $L$. The multi-objective partition allocation problem encompasses the following optimization objectives, the ultimate goal of which is to efficiently amplify the homogeneity of concatenated documents within the identical attention sequence:
@@
  • G1: Relevance. One fundamental goal is to maximize the holistic semantic coherence among documents contained in the same context window, with the observation that indicates a reduction in the perplexity of language models with increasing relevance (Shi et al., 2023).
  • G2: Homogeneity. Another objective to be pursued is diminishing cross-document attention during the model’s training phase, which can be accomplished by mitigating the quantity of documents utilized for concatenation within each sequence.
@@
  • G3: Integrity. Eliminating unnecessary truncation when concatenating documents for each training sequence is essential for preserving the integrity of individual documents (Ding et al., 2024), ensuring that the contextual richness and informational completeness are maintained to the greatest extent possible.
@@
  • G4: Compute Efficiency. In the process of managing vast training datasets for LLMs, efficiency emerges as a critical consideration. When aligning long-context training data with multi-objective goals, a balance among accuracy, processing speed, and space consumption is vital for enhancing LLM performance while optimizing resource efficiency.
@@
Soft constraints mandate that the sum of document length within each individual context window can not surpass the respective capacity, with the aim of maximizing document integrity and minimizing truncated token proportion within the same document (G3). Thus, Equation 1 can be written as
@@
where soft constraints are transformed into penalty terms, while $\alpha,\beta,\lambda$ are non-negative hyper-parameters that control the impact of objective and penalty functions.
The respective objective functions are
@@
Moreover, the penalty term $p(x)$ can be written as
(4) $p(x)=\begin{cases}0,\text{if }\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}\leq L&\\ \frac{L}{\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}-1,\text{otherwise}&\end{cases}$
which serves as a penalty for the truncation of concatenated documents that exceed context window capacity, thereby causing document incompleteness.
Solving MOCO problems exactly is NP hard, but coarse-to-fine algorithms are recognized for offering an efficient approximate solution (G4). However, the challenge lies in effectively reconciling multiple objects within the constraints of computational resources and the vast scale of the training corpus $\mathcal{D}$ (Zhao et al., 2005). We denote $|\mathcal{D}|$ as the number of documents contained in $\mathcal{D}$. The quantity of documents $k$ chosen for concatenation within a specific context is notably smaller than the total document count in $\mathcal{D}\,(k<<|\mathcal{D}|)$. Consequently, despite conducting score-based sorting on the remaining $|\mathcal{D}|-k$ documents during each iteration, they remain unselected, thereby consuming a significant expenditure of time and computational resources. Therefore, we initially perform semantic clustering on the entire training corpus $\mathcal{D}$ at a coarser granularity, followed by the implementation of a greedy algorithm to optimize the trade-off between efficiency and precision during the document organization process.
3.2. Phase 0: Data Preprocessing
Given the training context window length $L$ of the model, we tokenize raw documents in the training corpus $\mathcal{D}$ and divide them into segments, each comprising no more than $L$ tokens, which is dictated by the context length constraints. Subsequently, we utilize the embedding model to embed each segmented document $d_{i}$ to get $E\left(d_{i}\right)$.
@@ 3.3. Phase 1: Semantic Clustering
Input: Training corpus $\mathcal{D}$, Cluster number $N_{c}$ gained from Equation 6, Hyperparameters: distance threshold $\delta$, alteration threshold $\epsilon$, maximum iterations $T$
Parameter: Unclustered data $\mathcal{U}$, Alteration distance $dst$
Output: Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$
Define $\mu(c)$: center vector of cluster $c$
1:  Sample $N_{c}$ centroids in $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ randomly from $\mathcal{D}$, $\mu(c_{j})\leftarrow E(d_{j})\,\text{for all }j=1,...,N_{c}$
[…]
3:  for $t=1,2,\ldots,T$ do
4:     for each documnet $d_{i}$ in $\mathcal{D}$ do
5:        if $\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))>\delta$ or $t==T$ then
6:           Add $d_{i}$ to $c_{j^{*}}=\arg\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))$
7:        else
8:           $\mathcal{U}.append(d_{i})$
9:        end if
10:     end for
11:     for each $d_{u}$ in $\mathcal{U}$ do
12:        Add new cluster $c_{u}$, $\mu(c_{u})\leftarrow E(d_{u})$, $N_{c}\leftarrow N_{c}+1$
13:     end for
14:     for each cluster $c_{j}$ do
15:        $dst\leftarrow dst+||\mu(c_{j})-\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})||$
16:        Recalculate centroid $\mu(c_{j})\leftarrow\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})$
17:     end for
18:     for each pair of clusters $c_{i}$ and $c_{j}$ do
19:        if $cos(\mu(c_{i}),\mu(c_{j}))>\delta$ then
20:           Merge cluster $c_{i}$ and cluster $c_{j}$, $N_{c}\leftarrow N_{c}-1$
21:        end if
22:     end for
23:     if $dst<\epsilon$ then
24:        Break
25:     end if
26:     $dst\leftarrow 0$
27:  end for
28:  return  Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ where $\cup_{j=1}^{N_{c}}c_{j}=\mathcal{D}$
In the initial phase, the documents undergo a process of semantic clustering. To estimate the initial cluster number, we derive inspiration from the construction of a distance matrix by quantifying semantic distances between documents within the training corpus $\mathcal{D}$. We first divide $\mathcal{D}$ into $n$ subsets $D_{1},D_{2},...,D_{n}$, where each subset $D_{j}=\{d_{1},d_{2},...,d_{k}\}\subset\mathcal{D}$ and $\bigcup_{j=1}^{n}D_{j}=\mathcal{D}$. For each subset $D_{j}$, we compute the density $\rho_{j}$ of its distance matrix, which demonstrates the average dissimilarity score in relation to the entire subset.
@@
We first randomly select $N_{c}$ data samples from the entire training corpus $\mathcal{D}$ to serve as the initial clustering centers, which is based on the calculations from Equation 6. These sampled data points are then indexed using the FAISS vector searching library. Throughout the clustering process, for each document $d_{i}$, we employ the HNSW method (Malkov and Yashunin, 2018) from FAISS to retrieve its nearest cluster center $c_{j}$. If the semantic similarity between the current center and $d_{i}$ exceeds a predefined threshold $\delta$, then $d_{i}$ is assigned to cluster $c_{j}$. Conversely, if the similarity does not surpass $\delta$, $d_{i}$ is designated as a new cluster center. Moreover, as the iterative clustering progresses, adjustments to the cluster centers are made by calculating the mean of the document embedding vectors within each cluster, thereby establishing new cluster centroids. Additionally, if the similarity between any two clusters exceeds the threshold $\delta$, those clusters are merged. This approach not only refines the clustering process by dynamically adjusting cluster centers based on semantic coherence but also ensures that clusters remain distinct and meaningful by merging overly similar clusters, thereby enhancing the overall structure and interpretability of the clustering outcome.
@@ 3.4. Phase 2: Greedy Policy
Algorithm 2 Greedy Semantic-Driven Largest-Fit Allocation
@@
Input: Cluster $c_{j}\in\mathcal{C}$, Context sequence capacity $L$, Context sequence quantity $N_{w}$, Hyperparameters: impact of objective and penalty function $\alpha,\,\beta,\,\lambda$
@@
Parameter: Available context sequence $\mathcal{A}$, Semantic relevance $f_{1}$, Residual capacity proportion $f_{2}$, Truncation penalty $p$, Concatenation score $F$
Output: Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$
Define $r(s)$: remaining capacity of context sequence $s$
Define $l(d)$: token quantity of document $d$
Define $\mu(s)$: center vector of context sequence $s$
1:  $\left\{r(s_{l})\leftarrow L\right\}_{l=1}^{N_{w}}$: Initialize empty context sequences
2:  $S_{c_{j}}$: Sort $d_{i}\in c_{j}$ by token quantity in descending order
3:  for each documnet $d_{i}$ in $S_{c_{j}}$ do
4:     $\mathcal{A}=\left\{s_{l}|r(s_{l})\geq 0\right\}_{l=1}^{N_{w}}$
5:     if $\mathcal{A}=\emptyset$ then
6:        Break
7:     end if
8:     for each context sequence $s_{a}$ in $\mathcal{A}$ do
9:        $f_{1}(d_{i},s_{a})=cos(E(d_{i}),\mu(s_{a}))$
10:        $f_{2}(d_{i},s_{a})=r(s_{a})/L$
11:        $p(d_{i},s_{a})=\begin{cases}0,\text{if }l(d_{i})\leq r(s_{a})&\\ L/(L+l(d_{i})-r(s_{a})),\text{otherwise}&\end{cases}$
12:        $F(d_{i},s_{a})=\alpha f_{1}(d_{i},s_{a})+\beta f_{2}(d_{i},s_{a})+\lambda p(d_{i},s_{a})$
13:        Add $d_{i}$ to $s_{a^{*}}$ where $s_{a^{*}}=\arg\max_{s_{a}\in\mathcal{S}}F(d_{i},s_{a})$
14:     end for
15:     $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$
16:     $\mu(s_{a^{*}})\leftarrow((|s_{a^{*}}|-1)\mu(s_{a^{*}})+E(d_{i}))/|s_{a^{*}}|$
17:  end for
18:  return  Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$

### v2

@@ 2.3. Multi-Objective Combinatorial Optimization
where $x$ denotes a solution in the decision space $\chi\subset\mathbb{R}^{n}$, $\mathbf{f}:\chi\to\mathbb{R}^{m}$ represents an $m$-dimensional vector-valued objective function, and $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k}\,\text{where}\,k=1,...,n$ stands for a series of soft constraints. Importantly, the evaluation process is computationally expensive for each individual objective $f_{i}\left(x\right),i=1,...,m$. The complexity of MOCO challenges lies in the fact that improving the outcome of one objective often leads to the deterioration of another, making it necessary to find a balance or trade-off between them.
3. Observations and Challenges
3.1. Observations on Training Datasets
The training data is sourced from a variety of domains, such as web content, books, academic papers, and code repositories, with the composition and respective weight proportions shown on the right side of Figure 2. We conducted a statistical analysis of token count characteristics across various domains, as shown in Table 1. Additionally, we categorized tokenized documents into different length intervals based on the token counts of each data source. The proportional distributions across domains are visualized as a normalized stacked horizontal bar chart on the left size of Figure 2, offering a hierarchical view of the domains according to the prevalence of longer texts.
From Table 1 and Figure 2, we have observed that data from the book and academic paper domains predominantly consist of long documents, with over half exceeding 64K tokens in length. Academic papers, however, tend to fall within the medium-length range, with documents typically spanning 4K to 32K tokens. In contrast, web-sourced data is primarily composed of short documents, with the vast majority under 4K tokens. As a result, web-sourced data is not naturally suited for long-context training due to its shorter document lengths, which often requires the concatenation of multiple documents to form sequences of adequate length for long-context training.
Figure 3. Illustration of DataSculpt. In the data preprocessing stage, documents are initially divided into chunks with $\leq$ context length $L$, which are subsequently transformed into vector embeddings. During the semantic clustering phase, we implement a variant of the ISODATA algorithm, which is augmented with the FAISS vector searching library, to aggregate documents based on semantic similarity. Following this, a greedy semantic-driven largest-fit algorithm is utilized to arrange these clusters into sequences that are multi-objective optimally configured for long-context training.
3.2. Challenges
Given the token count statistics across documents, one of the primary challenges lies in the variation of sequence lengths among different data sources. In efforts to preserve domain diversity, shorter documents in the training dataset tend to increase the incidence of cross-document attention, where the model focuses on multiple, unrelated documents within the same context window. This can lead to model hallucinations and increased computational consumption. The problem becomes more prominent as the model’s context window expands during training.
Model Hallucinations. Model hallucinations occur when models generate incorrect or irrelevant content that is not grounded in the input data. While well-trained models generally reduce cross-document attention, this behavior is not always strictly controlled, which may potentially lead to spurious outputs as well as unintended correlations (Pouransari et al., 2024; Ding et al., 2024).
Computational Consumption. Moreover, cross-document attention expends unnecessary computational resources by focusing on unrelated tokens that contribute little to the model’s learning process (Krell et al., 2021; Variš and Bojar, 2021). The unnecessary processing of unrelated tokens not only diminishes the model’s performance but also amplifies the computational demands, which is especially crucial given the attention mechanism’s quadratic complexity (Vaswani et al., 2017).
3.3. Problem Formulation
Based on the above observations and analysis, we frame the scenario of data organization and management for long-context training as a multi-objective partition allocation problem with soft constraints at a massive data scale. This is fundamentally an instance of the multi-objective combinatorial optimization (MOCO) problem (Lin et al., 2022) described in Section 2.3.
@@ 3.3. Problem Formulation
Setup.  Given a training dataset $\mathcal{D}$ and training sequences $\mathcal{W}$ awaiting sampling, with each context window limited to a length of $L$, the multi-objective partition allocation problem focuses on optimizing several key objectives. The overarching goal is to efficiently enhance the homogeneity of concatenated documents within the same attention sequence:
@@
  • G1: Relevance. The primary goal is to maximize semantic coherence among documents within the same context window. Increasing relevance has been shown to reduce language model perplexity, improving overall model performance (Shi et al., 2023).
  • G2: Homogeneity. Reducing cross-document attention during the model’s training phase is crucial. This can be achieved by minimizing the number of concatenated documents within each context window, ensuring the model focuses on fewer, more cohesive sequences.
@@
  • G3: Integrity.Preserving the integrity of individual documents by minimizing unnecessary truncation during concatenation is essential. This ensures that the contextual richness and completeness of each document are retained as much as possible (Ding et al., 2024).
@@
  • G4: Efficiency.Efficiency is a key concern when managing large-scale training datasets for LLMs. Achieving a balance between accuracy, processing speed, and space consumption is critical for optimizing long-context training performance while minimizing resource usage.
@@
Soft constraints ensure that the total length of documents within each context window does not exceed its capacity, aiming to maximize document integrity and minimize the proportion of truncated tokens within the same document (G3). Consequently, Equation 2 can be written as:
@@
where soft constraints are incorporated as penalty terms, while $\alpha$, $\beta$, $\lambda$ are non-negative hyper-parameters that control the weight of each objective and penalty function.
The respective objective functions can be formulated as:
@@
Furthermore, the penalty term $p(x)$ is defined as:
(5) $p(x)=\begin{cases}1,\text{if }\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}l(d_{i})\leq L&\\ \frac{L}{\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}-1,\text{otherwise}&\end{cases}$
where $l(d_{i})$ represents the length of document $d_{i}$. The penalty term ensures that the total length of concatenated documents remains within the capacity $L$ of the context window. If the total sequence length exceeds $L$, the penalty increases to discourage overfitting the context window.
Solving MOCO problems exactly is NP-hard, but coarse-to-fine algorithms are known for providing efficient approximate solutions (G4). The primary challenge, however, lies in effectively balancing multiple objectives within the constraints of limited computational resources and the vast scale of the training corpus $\mathcal{D}$ (Zhao et al., 2005).
Let $|\mathcal{D}|$ denote the total number of documents in $\mathcal{D}$. The number of documents $k$ selected for concatenation within a given context window is significantly smaller than the total document count, such that $k<<|\mathcal{D}|$. As a result, even though score-based sorting is performed on the remaining $|\mathcal{D}|-k$ documents during each iteration, they remain unselected, consuming substantial time and computational resources.
To address this inefficiency, we first apply semantic clustering to the entire training corpus $\mathcal{D}$ at a coarser granularity. This is followed by a greedy algorithm that optimizes the trade-off between computational efficiency and precision during the document organization process.
4. DataSculpt
In this work, we introduce DataSculpt, a novel data management framework designed for long-context training. This framework aims to effectively integrate data from multiple domains and optimize the context window during training. Figure 3 presents the workflow of DataSculpt. In the data preprocessing phase, we first tokenize the raw documents from the training corpus, $\mathcal{D}$, according to the context window length, $L$, of the model. This process involves dividing the documents into segments, each containing no more than $L$ tokens, to comply with the model’s context length constraints. Following this, each document segment $d_{i}$ is processed through the embedding model to generate its embedding $E(d_{i})$.
@@ 4.1. Phase 1: Semantic Clustering
Input: Training corpus $\mathcal{D}$, Cluster number $N_{c}$ gained from Equation 7, Hyperparameters: distance threshold $\delta$, alteration threshold $\epsilon$, maximum iterations $T$;
Parameter: Unclustered data $\mathcal{U}$, Alteration distance $dst$;
Ouptut: Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$;
Define: $\mu(c)$ is center vector of cluster $c$;
1 Sample $N_{c}$ centroids in $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ randomly from $\mathcal{D}$, $\mu(c_{j})\leftarrow E(d_{j})\,\text{for all }j=1,...,N_{c}$ ;
[…]
3 for $t=1,2,\ldots,T$ do
4       for each documnet $d_{i}$ in $\mathcal{D}$ do
5             if $\max_{c_{j}\in\mathcal{C}}cos\left(E\left(d_{i}\right),\mu\left(c_{j}\right)\right)>\delta$ or $t==T$ then
6                   Add $d_{i}$ to $c_{j^{*}}=\arg\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))$;
7
8             $\mathcal{U}.append(d_{i})$;
9
10      for each $d_{u}$ in $\mathcal{U}$ do
11             Add new cluster $c_{u}$, $\mu(c_{u})\leftarrow E(d_{u})$, $N_{c}\leftarrow N_{c}+1$;
12
13      for each cluster $c_{j}$ do
14             $dst\leftarrow dst+||\mu(c_{j})-\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})||$;
15             Recalculate centroid $\mu(c_{j})\leftarrow\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})$;
16
17      for each pair of clusters $c_{i}$ and $c_{j}$ do
18             if $cos(\mu(c_{i}),\mu(c_{j}))>\delta$ then
19                   Merge cluster $c_{i}$ and cluster $c_{j}$, $N_{c}\leftarrow N_{c}-1$;
20
21      if $dst<\epsilon$ then
22             Break;
23
24      $dst\leftarrow 0$;
25
26return Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ where $\cup_{j=1}^{N_{c}}c_{j}=\mathcal{D}$ ;
@@ 24      $dst\leftarrow 0$;
In the initial phase, the documents undergo a process of semantic clustering at a relatively coarse granularity. To estimate the initial cluster number, we derive inspiration from the construction of a distance matrix by quantifying semantic distances between documents within the training corpus $\mathcal{D}$. We first divide $\mathcal{D}$ into $n$ subsets $D_{1},D_{2},...,D_{n}$, where each subset $D_{j}=\{d_{1},d_{2},...,d_{k}\}\subset\mathcal{D}$ and $\bigcup_{j=1}^{n}D_{j}=\mathcal{D}$. For each subset $D_{j}$, we compute the density $\rho_{j}$ of its distance matrix, which demonstrates the average dissimilarity score in relation to the entire subset.
@@
As shown in Algorithm 1, we first randomly select $N_{c}$ data samples from the entire training corpus $\mathcal{D}$ to serve as the initial clustering centers (Line 1), which is based on the calculations from Equation 7. These sampled data points are then indexed using the FAISS vector searching library. Throughout the clustering process with a maximum iteration count of $T$ (Line 3-18), for each document $d_{i}$, we employ the HNSW method (Malkov and Yashunin, 2018) from FAISS to retrieve its nearest cluster center $c_{j}$ (Line 4-7). If the semantic similarity between the current center and $d_{i}$ exceeds a predefined threshold $\delta$, then $d_{i}$ is assigned to cluster $c_{j}$ (Line 5-6). Conversely, if the similarity does not surpass $\delta$, $d_{i}$ is designated as a new cluster center (Line 7). Moreover, as the iterative clustering progresses, adjustments to the cluster centers are made by calculating the mean of the document embedding vectors within each cluster, thereby establishing new cluster centroids (Line 10-12). Additionally, if the similarity between any two clusters exceeds the threshold $\delta$, those clusters are merged (Line 13-15). This approach not only refines the clustering process by dynamically adjusting cluster centers based on semantic coherence but also ensures that clusters remain distinct and meaningful by merging overly similar clusters, thereby enhancing the overall structure and interpretability of the clustering outcome.
@@ 4.2. Phase 2: Greedy Policy
Input: Cluster $c_{j}\in\mathcal{C}$, Context sequence capacity $L$, Context sequence quantity $N_{w}$, Hyperparameters: impact of objective and penalty function $\alpha,\,\beta,\,\lambda$;
@@
Parameter: Available context sequence $\mathcal{A}$, Semantic relevance $f_{1}$, Residual capacity proportion $f_{2}$, Truncation penalty $p$, Concatenation score $F$;
Ouptut: Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$;
Define: $r(s)$ is remaining capacity of context sequence $s$;
Define: $l(d)$ is token quantity of document $d$;
Define: $\mu(s)$ is center vector of context sequence $s$;
1 $\left\{r(s_{l})\leftarrow L\right\}_{l=1}^{N_{w}}$: Initialize empty context sequences;
2 $S_{c_{j}}$: Sort $d_{i}\in c_{j}$ by token quantity in descending order;
3 for each documnet $d_{i}$ in $S_{c_{j}}$ do
4       $\mathcal{A}=\left\{s_{l}|r(s_{l})\geq 0\right\}_{l=1}^{N_{w}}$;
5       if $\mathcal{A}=\emptyset$ then
6             Break;
7
8      for each context sequence $s_{a}$ in $\mathcal{A}$ do
9             $f_{1}(d_{i},s_{a})=cos(E(d_{i}),\mu(s_{a}))$;
10             $f_{2}(d_{i},s_{a})=r(s_{a})/L$;
11             $p(d_{i},s_{a})=\begin{cases}1,\text{if }l(d_{i})\leq r(s_{a})&\\ L/(L+l(d_{i})-r(s_{a})),\text{otherwise}&\end{cases}$;
12             $F(d_{i},s_{a})=\alpha f_{1}(d_{i},s_{a})+\beta f_{2}(d_{i},s_{a})+\lambda p(d_{i},s_{a})$;
13             Add $d_{i}$ to $s_{a^{*}}$ where $s_{a^{*}}=\arg\max_{s_{a}\in\mathcal{S}}F(d_{i},s_{a})$;
14
15      $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$;
16       $\mu(s_{a^{*}})\leftarrow((|s_{a^{*}}|-1)\mu(s_{a^{*}})+E(d_{i}))/|s_{a^{*}}|$;
17
18return Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$;
Algorithm 2 Semantic-Driven Largest-Fit Allocation